sue rented a car, a one time deposit is required with the first month pay. she paid $950 the first montb abd a total of $9600 during the first year. what is the monthly rent and the amount of the deposit sue paid?
Answer by mananth(16949)

Let x be the deposit paid in the first month
9600-x is the rent for 12 months
rent for 1 month = (9600-x)/12 = 960-x


(9600-x)/12 = 960-x
9600=12(960-x)
9600-x =11520-12x
11x=1920
x=174.5
9600-174.54= 785.5
Deposit = 174.5
Monthly 785.5
m.ananth@hotmail.ca